Status Report

NASA Space Station On-Orbit Status 20 Feb 2004

By SpaceRef Editor
February 20, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 20 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

The station crew’s sleep cycle is now shifted forward again by two hours (wake: 11:00pm EDT last night; sleep: 3:00pm this afternoon).

Early in the morning, on Daily Orbit 3 over RGS (Russian ground sites), the crew conducted a review of the EVA-9 timeline for 2/26, supported by tagup with specialists at TsUP/Moscow via S-band.

After the timeline review, on DO 4, Kaleri and Foale prepared the SKK and Kromka payload equipment, to be taken outside on the spacewalk, followed by recording video and accompanying commentary on the DC-1 and equipment preps with ground specialists.    [Objectives of the spacewalk include installation of the Matryoshka radiation measuring equipment, photography of two Japanese contamination monitors (MPAC, SEED), removal and swapout of foldable sampling panels (SKKs, removable cassette containers), replacement of the Russian Kromka contamination monitoring SKK, ATV (automated transfer vehicle) support engineering for EVA (removal/relocation of LSV laser retroreflectors), and removal of a foreign object (cable strap) from the WA-2 ham radio antenna.  The crew today set up three sampling SKKs (#2, #3, #4) to replace the older SKKs mounted outside the DC-1 and Service Module (SM), and the Kromka SKK.  They also recorded video imagery of their work for the subsequent discussion with EVA specialists at TsUP.]

Both crewmembers in turn performed the CHeCS (crew health care systems) emergency medical operations OBT (on-board training) drill, a 30-min. exercise to refresh the Crew Medical Officer (CMO)’s acuity in applying ACLS (advanced cardio life support) in an emergency.   [Today’s computer-based proficiency drill focused on administration of intravenous (IV) fluid infusion and the diagnosis and treatment of a tension pneumothorax.  (The latter refers to a collection of gas in the pleural space in the chest resulting in collapse of the lung on the affected side.  A tension pneumothorax is a life-threatening condition, where the air within the pleural space is under pressure, displacing mediastinal structures and compromising cardiopulmonary function).]

After some more troubleshooting on the Elektron-VM oxygen generator by Sasha Kaleri, the ground succeeded in powering up the unit, and it has been running satisfactorily now for several hours on the primary pump.    [Background:  With a recently introduced new procedure to prevent gas bubbles from getting into the BZh fluid unit, the FE has to prime (fill) the KOV thermal loops’ EDV container with purified water from the multifiltration/purification column unit (BKO) while checking for any air bubbles. EDV water was not normally passed through the BKO, as is the case for water from a CWC (contingency water container).  The most recent series of Elektron offnominal shutdowns started on 2/10, and since then the unit has not been operating for longer than a 24-hour period.  On 2/13, its RPD pressure differential regulator was replaced to fix a noise problem that cropped up several weeks earlier and had gotten louder, but the noise was not completely eliminated, and air was/is still getting into the Elektron’s fluid unit (BZh).  On 2/18, voltage readings taken on the two micropumps were found nominal.  Yesterday, the crew performed a nitrogen (N2) purge of the BZh, intended to compress the buffer tank (BE) and force out any air bubbles trapped in the micropumps; the Elektron failed seven minutes after reactivation.  It is now up and running again.]

On the Vozdukh CO2 removal system, the FE performed the periodic functional closure test of its AVK emergency vacuum valves (last time done: 1/14).    [The AVKs are critical because they close the Vozdukh’s vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA).  Access to vacuum is required to vent carbon dioxide during the regeneration of the absorbent cartridges (PP).  During nominal operation, the AVK valves remain open.  Only two of the three absorbent channels are functioning.]

Preparatory for its subsequent calibration with the MCA (major constituents analyzer), Kaleri deactivated the IK0501 gas analyzer (GA) in the SM and exchanged its BF carbon dioxide (CO2) filter assembly with a new unit from the FGB.  The spent BF was stowed for disposal (replaced last: 1/9).  Sasha also adjusted the IK0501 for its subsequent calibration readings from its IG-3 O2 gas analyzer.   [IK0501 is an automated system for measuring CO2, O2, and H2O in the air, as well as the flow rate of the gas being analyzed.]

After the U.S. MAS was switched from life extending mode (LEM) to full operational mode by ground commands last night at 11:45pm EDT, the CDR opened its VGA (verification gas assembly) valve in support of a ground-controlled zero calibration of the analyzer six hours later.   [In LEM, the MCA is not sampling the ISS cabin air, but the ion pump is kept operating in order to maintain vacuum in the mass spectrometer, which allows the MCA to be returned easily and quickly to Operational status.  Its electrical current is greatly reduced in this mode, as is the rate of erosion of the ion pump cathodes, thus prolonging the pump’s life.  During the period in LEM, the MCA provides no atmosphere constituent partial pressure data.]

After the MCA zero calibration, Kaleri supported activation of the Russian IK0501 GA to take ppO2 measurements for calibrating the instrument with the MCA data obtained at about the same time.

Also in support of the instrument calibrations, Foale used the U.S. CDMK (carbon dioxide monitoring kit) to take the weekly reading of the cabin air’s CO2 partial pressure in the SM and Lab module, for calldown to the ground (along with the battery status) for use in trending analyses.

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer.

As every week, Mike transferred data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm.  Later, he completed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleting them on the HRM.  

In the Lab, the CDR completed the periodically required process of transferring the water collected in the Lab condensate tank to a CWC (using CWC #1042).  The offloading took about 30 minutes.  The necessary jumper/plumbing setup was then removed again.

Mike also conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO), and completed the regular routine status checkup of the autonomous Increment 8 payloads.

Sasha conducted his regular checkup of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which researches growth and development of plants under spaceflight conditions in the Lada-4 greenhouse.

Kaleri worked on the SM Central Post, replacing its Laptop 1 with Laptop 3, which up to then controlled some medical and payload experiments.  Moscow has requested U.S. assistance in repairing Laptop 1.   [Laptop 1, working with the Central Post computer (KTsP1), failed a week ago (2/12).  After operations were switched to the redundant (backup) laptop 2, with KTsP2, Kaleri replaced its HDD (hard disk drive) #6136 on 2/13 with a spare (#6059), but it again failed to reboot.  After today’s laptop swap, Laptop 3’s medical and payload experiments can no longer be operated, making repair of #1 relatively urgent, although #3 apparently is not required to support medical activities prior to or following the EVA-9.]

At 1:30pm, CDR Foale engaged in a ham radio pass with amateur radio fans at Glenwood Elementary School, Perrysburg, OH.   [The school is located in a suburb of Toledo, OH, and has an enrollment of 485 students in grades K through 6, of which a high percentage qualify for state and federally funded programs.  Students have participated in a videoconference with NASA JSC personnel, have welcomed speakers from NASA GRC and were visited by Astronaut Dr. Thomas.]

The crew had their weekly teleconference with the ISS Flight Director at MCC-H.

At dinnertime (12:45pm), as every day, the crew supported the Renal Stone prevention experiment by taking the test medication (either potassium citrate or placebo tablets) until the next sample collection phase in early April this year.

Shortly after the EVA/Orlan demo yesterday, MCC-H lost Ku- and S-band comm with the ISS for almost six minutes, when the required TDRS (Tracking & Data Relay Satellite) connection was not established as scheduled.  According to initial explanations from White Sands, the TDRS Spare did not perform its necessary pointing move for about five minutes, possibly due to a “hang-up” on the ground computer that provides real-time pointing commands.  The anomaly is under investigation.    [In the event of both Ku- and S-band loss, the ISS has only one alternate comm path, i.e., via VHF over Russian ground sites (or US ground sites if they have been activated).  A loss of this kind could severely impact any ongoing critical operations on ISS, such as an EVA.]

Moscow has prepared the reboost/phasing burn sequence required by Progress 13P to meet Soyuz 8S launch and 7S landing constraints.  Current dates are 3/2 (delta-V: 2.3 m/s); 3/31 (~2.5 m/s); and 4/10-16 (~2 m/s).  8S launch is scheduled on 4/19, 7S landing on 4/29.

The crew’s sleep time begins at 3:00pm EST.  Tonight they can sleep really long: until 5:45am tomorrow morning.

Today’s CEO targets, excluding almost all targets in the Americas due to dominantly descending orbit tracks combined with the sleep shift, were Seoul, South Korea (nadir pass), Taiwan Smog (very poor recent visibility over Taiwan due to smog began to improve in the last 12 hours.  Margins of the smog mass should have been visible in oblique views left and right of track as ISS passed over the southern part of the island), Ho Chi Minh City, Vietnam (just right of track.  Scattered cumulus cloud only was predicted.  Since this city of 5.5 million has a diffuse boundary, the ground suggested capturing a wider area than the overt cityscape), Mekong River Delta (detailed mapping swath requested to compare with prior images of land use in this rapidly changing region), Internal waves, S China Sea (looking right towards the glint disc for features generated over the shallow ocean bottom between Vietnam and Borneo.  The glint disc is relatively close to the spacecraft, which should allow more detailed views; and this is the optimum time of the month for internal waves to form), Lake Eyre, Australia (status of water levels requested for another data point in this long term monitoring site), Hyderabad, India (nadir pass over this major regional city of 3.7 million people and six universities), Addis Ababa, Ethiopia (a recent ISS/CEO image of this remote capital city is the best ground investigators have, but clouds masked part of the urban area so that another attempt was requested.  This is a nadir pass), and Saharan dust (Dynamic event.  ISS passed over the Saharan city of Agadez in Niger where visibility was reported reduced to 1 mile.  Looking left towards the dust source area in Chad where weather satellite images suggest thick blowing dust).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:34am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32A).  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.2; ppO2 (mmHg) — 159.4; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 21.4.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 23.7.
  • Node:  Pressure (mmHg) — 739.44; temperature (deg C) — 23.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 742.59; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 742.80; temperature (deg C) — 23.7; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.4
  • PMA-2:  Shell heater temp (deg C) — 12.9

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #6 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 5:52am EST [= epoch]):

  • Mean altitude — 366.6 km
  • Apogee — 371.1km
  • Perigee — 362.1 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0006683
  • Orbits per 24-hr. day — 15.67
  • Solar Beta Angle — -46.1 deg (magnitude decreasing)
  • Mean altitude loss last 24 hours — 90 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29984

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.